# Why We Build Local Large Language Models: An Observational Analysis from 35 Japanese and Multilingual LLMs

### v1

@@ 4.2 Correlation Between Evaluation Benchmarks and Language-Specific Performance
To group benchmarks based on the similarities of LLM performance, we computed a Pearson correlation between two benchmarks $a$ and $b$. More specifically, let the column vectors $X_{:,a}$ and $X_{:,b}$ represent the array of two benchmarks $a$ and $b$, we compute the Pearson correlation coefficient $\mathrm{corr}(X_{:,a},X_{:,b})$. Figure 3 shows the Pearson correlation matrix, revealing two key findings [^7] We confirmed that using Spearman’s rank correlation produced no significant differences in the findings.:
@@
Second, QA tasks about Japanese knowledge (JEMHopQA, NIILC) and an English-Japanese translation task (WMT20-en-ja) exhibit relatively weak correlations with other tasks respectively. In particular, NIILC shows negative correlations with most English tasks, and WMT20-en-ja does nearly no correlations with them. These facts suggest that performance on these tasks may be determined by factors different from those influencing other tasks.
While we observe strong linear correlations between JMMLU, MGSM, and JHumanEval and their English counterparts, given that these are derived from English benchmarks, readers may be concerned that cross-lingual correlations of these benchmarks are overestimated. An straightforward workaround would be to evaluate using random, non-overlapping subsets of instances for each language. Instead of implementing this directly, we approximated the accuracy variation introduced by such splits using the estimated standard error (SE) following Biderman et al. (2024) and confirmed that impact of the fluctuation by the SE is negligible on the observed linear trends. For example, MGSM has 250 instances, and the SE for an accuracy of 0.5 is approximately $\sqrt{\frac{0.5(1-0.5)}{250}}\approx 0.032$. In contrast, the observed standard deviation of accuracy across LLMs was 0.246, sufficiently larger than the SE.
@@ 4.3 Principal Component Analysis (PCA)
Formally, we first standardize each column of $X$ to have a mean of zero and a standard deviation of one: $\hat{X}$. Next, we perform eigendecomposition of the correlation matrix as $\hat{X}^{\top}\hat{X}=U\Lambda U^{\top}$, where $U=[u_{1},u_{2},\dots,u_{D}]$, and $u_{j}\in\mathbb{R}^{D}$ is the $j$-th unit-length eigenvector. We then select the top four principal components (PCs), as their cumulative fraction of variance explained ($r$; contribution ratio) is 90.8% (= 65.2% + 15.4% + 7.0% + 3.2% from PC1 to PC4). We define the eigenvectors corresponding to PC1 to PC4, $U_{4}=[u_{1},u_{2},u_{3},u_{4}]\in\mathbb{R}^{D\times 4}$ as the factor loadings and compute corresponding PC scores as $S_{4}=\hat{X}U_{4}$. Given that $U$ is an orthonormal matrix and the total variance explained by PC1–PC4 is approximately 90%, the original matrix can be approximated as the product of PC scores and factor loadings: $\hat{X}\approx S_{4}U_{4}^{\top}$.
@@
In this way, we decompose standardized benchmark scores $\hat{X}$ into the product of LLM-specific principal component scores (ability factors) $S_{4}\in\mathbb{R}^{M\times 4}$ in Figure 3 and benchmark-specific factor loadings $U_{4}\in\mathbb{R}^{D\times 4}$ in Figure 4, which represent the associations between the ability factors and task performances [^9] Technically, since the signs and magnitudes of the PC scores and factor loadings are arbitrary, we adjusted the signs for ease of interpretation and normalized the factor loading vectors to have an $L_{2}$ norm of 1..
@@
The first principal component (PC1) has relatively uniform factor loadings. As shown in Figure 5 left, LLMs with higher PC1 scores tend to have higher average benchmark scores in both English and Japanese, suggesting that PC1 represents a general ability factor. It represents the average performance across most benchmark scores, including commonsense and reading comprehension in Japanese. It is also noteworthy that the factor loadings for the three Japanese benchmarks, NIILC, WMT20-en-ja, and JEMHopQA, are relatively small, suggesting that these benchmark scores are more closely associated with a different principal component.
@@
The second principal component (PC2) shows concentrated factor loadings on JEMHopQA, NIILC, and WMT20-en-ja, and relatively small factor loadings on JCommonsenseQA and JSQuAD, indicating the abilities of (encyclopedic) knowledge about Japan and English-Japanese translation. In fact, Figure 3 shows that LLMs pre-trained on Japanese text, such as Swallow and Sarashina2 families, have high PC2 scores, which will be analyzed in detail in § 4.4. Additionally, as shown in Figure 5 center, the higher PC2, the higher benchmark scores on those tasks. For instance, the magin of NIILC accuracy between LLMs with the lowest and highest PC2 scores is approximately 40 points. Considering that PC1 has relatively low factor loadings for these benchmarks, PC2 represents Japanese-specific abilities, such as QA about Japanese knowledge and English-Japanese translation, whereas PC1 represents the general abilities.
@@
The third principal component (PC3) shows concentrated factor loadings on MGSM, GSM8K, JHumanEval, and HumanEval, representing abilities of multilingualism, language-agnostic arithmetic reasoning, and code generation. As shown in Figure 5 right, there is a moderate trend suggesting that higher PC3 score are associated with higher benchmark scores on code-generation and arithmetic-reasoning. While we observe some outliers in the lower right corner, which correspond to LLM-jp-13B v2.0, CyberAgentLM2-7B, and Fugaku-LLM 13B, we think the PC3 scores for these LLMs might be overestimated due to compensation for their excessively low PC1 scores (as seen in Figure 3). This interpretation is supported by the alternative factor analysis using Promax rotation (Appendix C.1), where we observed diminished principal component scores for arithmetic reasoning and code generation in these LLMs.
@@ 4.4 Scaling Laws between Ability Factors and Computational Budget
Figure 6 shows the scatter plot with the English computational budget (log scale) and PC1. It reveals that the general ability (PC1) scales with the English computational budget (Pearson’s $\rho=0.916$) [^10] The correlation with the logarithm of the total computational budget was slightly higher ($\rho=0.938$). Still, given the weak correlation with the Japanese computational budget, we concluded that it scales more with the English computational budget.. Figure 7 shows the scatter plot with the Japanese computational budget (log scale) and PC2. We can see that the Japanese ability (PC2) moderately scales with the Japanese computational budget ($\rho=0.779$). We also confirmed that the correlation between PC2 and the English or total computational budget is much weaker ($\rho=0.164$ and $0.186$, respectively). These findings indicate that PC2 and associated Japanese task performances scale with an increase in Japanese training tokens, thereby supporting our claim in § 4.3 that “PC2 represents Japanese ability.” Furthermore, we argue that the source of Japanese ability lies in the computational budget allocated to Japanese texts.
4.5 PCA for Japanese LLMs Trained from Scratch
In order to remove the influence of high compute budgets for English LLMs, we excluded LLMs continually pre-trained in Japanese and focused only on 20 LLMs trained from scratch. Figure 8 shows factor loadings of PCs with the 20 LLMs, which also identified similar ability factors to those found in § 4.3. We omit the results of relationships between computational budgets and English and Japanese abilities, but observed the consistent correlations with Figures 6 and 7 (see Figures 13 and 14 in Appendix C.2).
@@ 5 Conclusion and Future Work
In this paper, we evaluated the performance of 35 Japanese, English, and Multilingual LLMs on 19 task benchmarks that assess the abilities in both Japanese and English. We then analyzed the cross-task and cross-lingual correlations of benchmark scores, mapped the performance in a low-dimensional ability space, and explored the relationship between ability factors and computational budgets for English and Japanese. The correlation analysis showed strong multilingual abilities in academic knowledge, code generation, and arithmetic reasoning tasks. This suggests that, in order to enhance the abilities of these tasks, there is no strong motivation for using Japanese training data.
@@
From these analyses, we concluded that the advantage of building local LLMs by training on Japanese text is particularly evident in acquiring local knowledge written in Japanese and enhancing the ability to translate from English. This conclusion is likely to characterize Japanese LLMs.
We consider two directions as future work. First, we plan to extend the analysis with more LLMs and evaluation tasks to discover additional insights. This includes using LLMs with unique designs, for example, Phi family (Li et al., 2023; Abdin et al., 2024), which were trained on synthetic text. We also want to add evaluation tasks such as Japanese logical reasoning and standardized admission exams. The second direction is to extend our analysis and findings to other languages. We believe that the conclusion of this paper can be generalized to: the advantage of building local LLMs by training in a language is acquiring local knowledge written in the language and enhancing the ability to translate from English to the language. This direction is nontrivial because conducting LLM experiments properly requires a deep understanding of the target languages and cultures. We hope this paper serves as a catalyst for analysis in other languages.
Limitations
Observational Approach
This study uses an observational approach, relying solely on the evaluation results of the LLMs. We did our best to collect open LLMs and benchmarks that are available as of writing and to evaluate LLMs correctly by ourselves. Still, the findings, including those from factor analysis, may be influenced by the selection of models and evaluation task benchmarks. Although we assessed the statistical error of factor loadings using leave-one-out cross-validation on the analyzed LLMs (see Figure 17 in the appendix) and confirmed that the standard deviations were small relative to the absolute values, this does not guarantee that our findings remain valid as a new best practice for designing LLMs emerges.
English Predominance in the Training Data
In § 4.4, we discussed that the general ability was acquired through training on English text, based on the solid log-linear relationship between the computational budget for English and PC1. However, this could be a limitation of our observational approach. Specifically, most of the LLMs evaluated in this study were trained on the English-centric data, with at least half of the data in English. When we have LLMs predominantly trained in Japanese, different findings might emerge.
Generalization of Findings beyond Japanese
This study focuses solely on Japanese LLMs as an instance of non-English and local LLMs. It is unclear whether the findings are applicable to other non-English LLMs. However, we want to emphasize that even evaluating LLMs accurately in languages that we are familiar with is not an easy task. Some LLMs scored zero for a benchmark, and we ended up debugging the problem only to find that they require a special token or even a line break in the prompt to obtain a valid generation. Implementation details (e.g., probabilistic decoding) also affected the performance of LLMs in downstream tasks. Technology has not yet advanced to the point where simply submitting a local LLM to a leaderboard yields reliable evaluation results effortlessly. Therefore, accurately conducting evaluation experiments with a lot of LLMs and benchmarks, as done in this study, requires a deep understanding of the target language, which sets a high bar for us.
Ethical Considerations
@@
Acknowledgments
This paper is based on results obtained from AIST policy-based budget project "R&D on Generative AI Foundation Models for the Physical Domain".
In the evaluation experiments of the LLMs, we utilized data and insights developed and made publicly available by LLM-jp.
Appendix A Details of the Experimental Setup
A.1 Evaluated Models
Table 1 shows a list of LLMs evaluated in this study. The table includes the name, the number of active parameters during inference, the base model from which the model was continually pre-trained, the language distribution of the training corpus, the total number of training tokens, the reported or estimated number of training tokens in English and Japanese, and the reference of each model. § A.3 explains the method used to estimate the number of language-specific training tokens. CPT stands for continual pre-training.
@@ A.1 Evaluated Models
Model name Num of params Con- stru- ction met- hod Source of CPT Corpus Training tokens EN tokens JA tokens Reference
Yi-1.5 6B 6 PT ZH,EN, Code 3600 2170 AI et al. (2024)
CyberAgentLM2-7B 7 PT JA,EN 1300 650 650 cyberagent/calm2-7b
Japanese Stable LM Base Gamma 7B 7 CPT Mistral-7B-v0.1 JA,EN 100 stabilityai/japanese-stablelm-base-gamma-7b
Japanese StableLM Beta 7B 7 CPT Llama2 7B JA,EN 2100 1794 102 stabilityai/japanese-stablelm-base-beta-7b
Llama 2 7B 7 PT EN 2000 1794 2 Touvron et al. (2023)
Mistral-7B-v0.1 7 PT EN Jiang et al. (2023)
Mistral-7B-v0.2 7 PT EN Jiang et al. (2023)
Qwen1.5-7B 7 PT 4000 2000 Team (2024)
Qwen2-7B 7 PT ZH,EN, Code+27 7000 3500 Yang et al. (2024)
RakutenAI-7B 7 CPT Mistral-7B-v0.1 JA,EN 175 RakutenGroup et al. (2024)
Sarashina2-7B 7 PT JA,EN 2100 840 1050 sbintuitions/sarashina2-7b
Swallow 7B 7 CPT Llama2 7B JA,EN 2100 1794 102 Fujii et al. (2024)
Swallow-MS v0.1 7 CPT Mistral-7B-v0.1 JA,EN, Code 100 Fujii et al. (2024)
Youri 7B 7 CPT Llama2 7B JA,EN 2040 1834 42 Sawada et al. (2024)
Llama 3 8B 8 PT EN 15000 14250 15 Dubey et al. (2024)
Llama 3 Swallow 8B 8 CPT Llama3 8B JA,EN, Code 15100 14250 115 Fujii et al. (2024)
Llama 3 Youko 8B 8 CPT Llama3 8B JA,EN 15022 14250 37 Sawada et al. (2024)
Yi-1.5 9B 9 PT ZH,EN, Code 3100 2170 AI et al. (2024)
ELYZA-japanese- Llama-2-13b 13 CPT Llama2 13B JA 2018 1794 20 Sasaki et al. (2023)
Fugaku-LLM 13B 13 PT JA,EN 400 200 200 Fugaku-LLM/Fugaku-LLM-13B
Llama 2 13B 13 PT EN 2000 1794 2 Touvron et al. (2023)
LLM-jp-13B v2.0 13 PT JA,EN, Code 260 120 130 LLM-jp et al. (2024)
Sarashina2-13B 13 PT JA,EN 2100 840 1050 sbintuitions/sarashina2-13b
Swallow 13B 13 CPT Llama2 13B JA,EN 2100 1794 102 Fujii et al. (2024)
Yi-1.5 34B 34 PT ZH,EN, Code 3100 2170 AI et al. (2024)
C4AI Command- R v0.1 35 PT JA,EN, ZH+8 CohereForAI/c4ai-command-r-v01
Mixtral-8x7B- v0.1 12.879 PT EN Jiang et al. (2024)
Swallow-MX 8x7B v0.1 12.879 CPT Mixtral-8x7B- Instruct-v0.1 JA,EN 100 Fujii et al. (2024)
Japanese Stable LM Beta 70B 70 CPT Llama2 70B JA,EN 2100 1794 102 stabilityai/japanese-stablelm-base-beta-70b
KARAKURI LM 70B v0.1 70 CPT Llama2 70B JA,EN 2016 1794 18 KARAKURI Inc. (2024)
Llama 2 70B 70 PT EN 2000 1794 2 Touvron et al. (2023)
Llama 3 70B 70 PT EN 15000 14250 15 Dubey et al. (2024)
Llama 3 Swallow 70B 70 CPT Llama3 70B JA,EN, Code 15100 14250 115 Fujii et al. (2024)
Swallow 70B 70 CPT Llama2 70B JA,EN 2100 1794 102 Fujii et al. (2024)
Qwen2-72B 72 PT ZH,EN, Code+27 7000 3500 Yang et al. (2024)
A.2 Evaluation Tasks and Benchmarks

### v2

@@ 4.2 Correlation Between Evaluation Benchmarks and Language-Specific Performance
To group benchmarks based on the similarities of LLM performance, we computed a Pearson correlation between two benchmarks $a$ and $b$. More specifically, let the column vectors $X_{:,a}$ and $X_{:,b}$ represent the array of two benchmarks $a$ and $b$, we compute the Pearson correlation $\mathrm{corr}(X_{:,a},X_{:,b})$. Figure 3 shows the Pearson correlation matrix, revealing two key findings [^6] We confirmed that using Spearman’s rank correlation produced no significant differences in the findings.:
@@
Second, QA tasks about Japanese knowledge (JEMHopQA, NIILC) and an English-Japanese translation task (WMT20-en-ja) exhibit relatively weak correlations with other tasks respectively. In particular, NIILC shows negative correlations with most English tasks, and WMT20-en-ja shows almost no correlations with them. These facts suggest that performance on these tasks may be determined by factors different from those influencing other tasks.
While we observe strong linear correlations between JMMLU, MGSM, and JHumanEval and their English counterparts, given that these are derived from English benchmarks, readers may be concerned that cross-lingual correlations of these benchmarks are overestimated. A straightforward workaround would be to evaluate using random, non-overlapping subsets of instances for each language. Instead of implementing this directly, we approximated the accuracy variation from random splits using the estimated standard error (SE) following Biderman et al. (2024) and confirmed that impact of fluctuation by the SE is negligible on the observed linear trends. For example, MGSM has 250 instances, and the SE for an accuracy of 0.5 is approximately $\sqrt{0.5(1-0.5)/250}\approx 0.032$. In contrast, the observed standard deviation of accuracy across LLMs was 0.246, sufficiently larger than the SE.
@@ 4.3 Principal Component Analysis (PCA)
Formally, we first standardize each column of $X$ to have mean of zero and a standard deviation of one: $\hat{X}$. Next, we perform eigendecomposition of the correlation matrix as $\hat{X}^{\top}\hat{X}=U\Lambda U^{\top}$, where $U=[u_{1},u_{2},\dots,u_{D}]$, and $u_{j}\in\mathbb{R}^{D}$ is the $j$-th unit-length eigenvector. We then select the top four principal components (PCs), as their cumulative variance explained ($r$; contribution ratio) is 90.8% (= 65.2% + 15.4% + 7.0% + 3.2% from PC1 to PC4). We define the eigenvectors corresponding to PC1 to PC4, $U_{4}=[u_{1},u_{2},u_{3},u_{4}]\in\mathbb{R}^{D\times 4}$ as the factor loadings and compute corresponding PC scores as $S_{4}=\hat{X}U_{4}$. Given that $U$ is an orthonormal matrix and the total variance explained by PC1–PC4 is about 90%, the original matrix can be approximated as the product of PC scores and factor loadings: $\hat{X}\approx S_{4}U_{4}^{\top}$.
@@
In this way, we decompose standardized benchmark scores $\hat{X}$ into the product of LLM-specific principal component scores (ability factors) $S_{4}\in\mathbb{R}^{M\times 4}$ in Figure 3 and benchmark-specific factor loadings $U_{4}\in\mathbb{R}^{D\times 4}$ in Figure 4, which represent the associations between the ability factors and task performances [^8] Since the signs and magnitudes of the PC scores and factor loadings are arbitrary, we adjusted the signs for ease of interpretation and normalized the factor loading vectors to have an $L_{2}$ norm of 1..
@@
The first principal component (PC1) has relatively uniform factor loadings. As shown in Figure 5 left, LLMs with higher PC1 scores tend to have higher average benchmark scores in both English and Japanese, suggesting that PC1 represents a general ability factor. It represents the average performance across most benchmark scores, including commonsense and reading comprehension in Japanese. This indicates that, unlike prior studies (§ 2.1), training on English text is effective and that Japanese-specific training is not necessarily for improving these abilities.
@@
The second principal component (PC2) shows concentrated factor loadings on JEMHopQA, NIILC, and WMT20-en-ja, and relatively small factor loadings on JCommonsenseQA and JSQuAD, indicating the abilities of (encyclopedic) knowledge about Japan and English-Japanese translation. In fact, Figure 3 shows that LLMs pre-trained on Japanese text, such as Swallow and Sarashina2 families, have high PC2 scores, which will be analyzed in detail in § 4.4. Additionally, as shown in Figure 5 center, the higher PC2, the higher benchmark scores on those tasks. For instance, the magin of NIILC accuracy between LLMs with the lowest and highest PC2 scores is approximately 40 points. Considering that PC1 has relatively low factor loadings for these benchmarks, PC2 represents Japanese-specific abilities, such as QA about Japanese knowledge and English-Japanese translation. Given that PC2 strongly associates with Japanese knowledge-based QA tasks, this aligns with previous work (Romanou et al., 2025), which found that multilingual LLMs struggle with cultural questions, especially in languages not included in the pre-training data.
@@
The third principal component (PC3) shows concentrated factor loadings on MGSM, GSM8K, JHumanEval, and HumanEval, representing abilities of multilingualism, language-agnostic arithmetic reasoning, and code generation. As shown in Figure 5 right, there is a moderate trend suggesting that higher PC3 score are associated with higher benchmark scores on code-generation and arithmetic-reasoning.
@@ 4.4 Scaling Laws between Ability Factors and Computational Budget
Figure 7 shows the scatter plot with the English computational budget (log scale) and PC1. It reveals that the general ability (PC1) scales with the English computational budget (Pearson’s $\rho=0.916$) [^9] The correlation with the logarithm of the total computational budget was slightly higher ($\rho=0.938$). Still, given the weak correlation with the Japanese computational budget, we concluded that it scales more with the English computational budget.
Figure 7 shows the scatter plot with the Japanese computational budget (log scale) and PC2. We can see that the Japanese ability (PC2) moderately scales with the Japanese computational budget ($\rho=0.779$). We also confirmed that the correlation between PC2 and the English or total computational budget is much weaker ($\rho=0.164$ and $0.186$, respectively). These findings indicate that PC2 and associated Japanese task performances scale with an increase in Japanese training tokens, thereby supporting our claim in § 4.3 that “PC2 represents Japanese ability.” Furthermore, we argue that the source of Japanese ability lies in the computational budget allocated to Japanese texts.
4.5 PCA for LLMs Trained from Scratch
To verify that our findings are not heavily influenced by the pre-training method, we repeated the analysis after excluding continually pre-trained Japanese LLMs, retaining only 20 LLMs trained from scratch. Figure 8 shows the factor loadings of PCs extracted from the performance of these 20 LLMs, revealing ability factors similar to those identified in the original analysis (§ 4.3). We omit the results of relationships between computational budgets and English and Japanese abilities, but observed the consistent correlations with Figures 7 and 7 (see Figures 14 and 14 in Appendix D.2).
@@ 5 Conclusion and Future Work
In this paper, we performed the most comprehensive evaluation to date, testing 35 Japanese, English, and Multilingual LLMs on 19 task benchmarks that assess the abilities in both Japanese and English. This breadth of coverage is one of the key novelties of our study and enables us to extract more generalizable insights than prior work. We then analyzed the cross-task and cross-lingual correlations of benchmark scores, mapped the performance in a low-dimensional ability space, and explored the relationship between ability factors and computational budgets for English and Japanese. The correlation analysis showed strong multilingual abilities in academic subjects, code generation, and arithmetic reasoning tasks. This suggests that, in order to enhance the abilities of these tasks, there is no strong motivation for using Japanese training data.
@@
From these analyses, we concluded that the advantage of building local LLMs by training on Japanese text is particularly evident in acquiring local knowledge written in Japanese and enhancing the ability to translate from English. This conclusion is likely to characterize Japanese LLMs. Our study is the first broad, unified evaluation across dozens of LLMs and an extensive benchmark suite to reveal which tasks do and do not benefit from target-language training.
We consider two directions as future work. First, we plan to extend the analysis with more LLMs and evaluation tasks to discover additional insights. This includes using LLMs with unique designs, for example, Phi family (Li et al., 2023; Abdin et al., 2024), which were trained on synthetic text. We also want to add evaluation tasks such as Japanese logical reasoning and standardized admission exams. The second direction is to extend our analysis and findings to other languages. We believe that the conclusion of this paper can be generalized to: the advantage of building local LLMs by training in a language is acquiring local knowledge written in the language and enhancing the ability to translate from English to the language. This direction is nontrivial because conducting LLM experiments properly requires a deep understanding of the target languages and cultures. We hope this paper serves as a catalyst for the development and anlaysis of non-English LLMs.
Acknowledgments
This work was supported by the ABCI Large-scale Language Model Building Support Program of the AI Bridging Cloud Infrastructure (ABCI), built and operated by the National Institute of Advanced Industrial Science and Technology (AIST).
Ethics Statement
@@ Acknowledgments
Reproducibility Statement
[…]
Broader Impacts
We believe our findings will contribute to the development of non-English LLMs. Moreover, this could foster a society in which every country has access to LLMs specialized in its own language and knowledge, thereby reducing the digital divide.
Appendix A Survey of Prior Work and Comparison with Our Analysis
We systematically surveyed prior works on non-English LLM development in two perspectives: coverage of design choices and effects of training on target languages.
[…]
More importantly, we observed the contradictory evidence for “language-agnostic” tasks. The majority of prior studies actually report gains from target-language training on commonsense reasoning (8 positive, 1 neutral, 0 negative) and academic subject benchmarks (5 positive, 2 neutral, 0 negative). These findings contrast both with our results. Furthermore, there seems no clear consensus on other tasks. For reading comprehension and mathematical reasoning benchmarks, prior work offers mixed or inconclusive evidence regarding the impact of target-language data (6 positive, 3 neutral, 0 negative; 3 positive, 2 neutral, 1 negative, respectively).
Table 1: The impact of training on the target language text. $\nearrow$: Positive, $\searrow$: Negative, $\rightarrow$: Neutral, -: Not investigated
Reference Lang Method
Read-
Compr.
Com-
mon-
Sense
Reason.
Math-
Reason.
MT to
Tgt Lang
MT from
Tgt Lang
Acad.
Subject
Coding
Local
Knowl.
QA
Sum-
mar-
iza-
tion
Ours JA
PT
CPT
$\rightarrow$ $\rightarrow$ $\rightarrow$ $\nearrow$ $\nearrow$ $\rightarrow$ $\rightarrow$ $\nearrow$ $\nearrow$
Etxaniz et al. (2024b) EU CPT $\nearrow$ $\nearrow$ $\nearrow,\rightarrow$ - - $\nearrow$ - $\nearrow$ -
Fujii et al. (2024) JA CPT $\nearrow$ $\nearrow$ $\nearrow$ $\nearrow$ $\searrow$ - - $\nearrow$ $\searrow$
Phasook et al. (2024) TH CPT $\rightarrow$ $\nearrow$ $\nearrow$ - - $\nearrow,\rightarrow$ - - -
Dou et al. (2024)
VI
TH
ID
MS
LO
CPT $\nearrow$ $\nearrow$ - - - $\nearrow,\rightarrow$ - - -
Joshi et al. (2025) HI CPT $\nearrow$ $\nearrow$ - - - $\nearrow$ - $\nearrow$ $\nearrow$
Vo et al. (2024) KO CPT $\nearrow$ $\nearrow$ - - - - - - -
Choi et al. (2024) KO CPT $\rightarrow$ $\nearrow$ - - - - - - -
Toraman (2024) TR CPT $\rightarrow$ $\rightarrow$ - - - - - - -
Larcher et al. (2023) PT CPT $\nearrow$ - - - - - - - -
Tejaswi et al. (2024b)
TA
HI
AR
TR
CPT - $\nearrow$ - $\nearrow$ - - - - $\nearrow$
Cui et al. (2024) ZH CPT - - - - - $\nearrow$ - - -
Etxaniz et al. (2024a) EU CPT - - - - - - - $\nearrow$ -
Holmström et al. (2023) SV PT - - $\searrow,\rightarrow$ - - - - - -
Luukkonen et al. (2023) FI CPT - - - - - - - - -
Pipatanakul et al. (2023) TH CPT - - - - - - - - -
Appendix B Details of the Experimental Setup
B.1 Evaluated Models
Table 2 shows a list of LLMs evaluated in this study. The table includes the name, the number of active parameters during inference, the base model from which the model was continually pre-trained, the language distribution of the training corpus, the total number of training tokens, the reported or estimated number of training tokens in English and Japanese, and the reference of each model. § B.3 explains the method used to estimate the number of language-specific training tokens. CPT stands for continual pre-training.
@@ B.1 Evaluated Models
Model name
Num
of
params
Con-
stru-
ction
met-
hod
Source of CPT
Corpus
Training
tokens
EN
tokens
JA
tokens
Reference
Yi-1.5 6B
6
PT
ZH,EN,
Code
3600 2170 AI et al. (2024)
CyberAgentLM2-7B
7
PT
JA,EN
1300 650 650 cyberagent/calm2-7b
Japanese Stable LM
Base Gamma 7B
7
CPT
Mistral-7B-v0.1
JA,EN
100 stabilityai/japanese-stablelm-base-gamma-7b
Japanese StableLM
Beta 7B
7
CPT
Llama2 7B
JA,EN
2100 1794 102 stabilityai/japanese-stablelm-base-beta-7b
Llama 2 7B
7
PT
EN
2000 1794 2 Touvron et al. (2023)
Mistral-7B-v0.1
7
PT
EN
Jiang et al. (2023)
Mistral-7B-v0.2
7
PT
EN
Jiang et al. (2023)
Qwen1.5-7B
7
PT
4000 2000 Team (2024)
Qwen2-7B
7
PT
ZH,EN,
Code+27
7000 3500 Yang et al. (2024)
RakutenAI-7B
7
CPT
Mistral-7B-v0.1
JA,EN
175 RakutenGroup et al. (2024)
Sarashina2-7B
7
PT
JA,EN
2100 840 1050 sbintuitions/sarashina2-7b
Swallow 7B
7
CPT
Llama2 7B
JA,EN
2100 1794 102 Fujii et al. (2024)
Swallow-MS v0.1
7
CPT
Mistral-7B-v0.1
JA,EN,
Code
100 Fujii et al. (2024)
Youri 7B
7
CPT
Llama2 7B
JA,EN
2040 1834 42 Sawada et al. (2024)
Llama 3 8B
8
PT
EN
15000 14250 15 Dubey et al. (2024)
Llama 3 Swallow 8B
8
CPT
Llama3 8B
JA,EN,
Code
15100 14250 115 Fujii et al. (2024)
Llama 3 Youko 8B
8
CPT
Llama3 8B
JA,EN
15022 14250 37 Sawada et al. (2024)
Yi-1.5 9B
9
PT
ZH,EN,
Code
3100 2170 AI et al. (2024)
ELYZA-japanese-
Llama-2-13b
13
CPT
Llama2 13B
JA
2018 1794 20 Sasaki et al. (2023)
Fugaku-LLM 13B
13
PT
JA,EN
400 200 200 Fugaku-LLM/Fugaku-LLM-13B
Llama 2 13B
13
PT
EN
2000 1794 2 Touvron et al. (2023)
LLM-jp-13B v2.0
13
PT
JA,EN,
Code
260 120 130 LLM-jp et al. (2024)
Sarashina2-13B
13
PT
JA,EN
2100 840 1050 sbintuitions/sarashina2-13b
Swallow 13B
13
CPT
Llama2 13B
JA,EN
2100 1794 102 Fujii et al. (2024)
Yi-1.5 34B
34
PT
ZH,EN,
Code
3100 2170 AI et al. (2024)
C4AI Command-
R v0.1
35
PT
JA,EN,
ZH+8
CohereForAI/c4ai-command-r-v01
Mixtral-8x7B-
v0.1
13101010Number of active parameters on inference. The total number of parameters is 47B.
PT
EN
Jiang et al. (2024)
Swallow-MX 8x7B
v0.1
13footnotemark:
CPT
Mixtral-8x7B-
Instruct-v0.1
JA,EN
100 Fujii et al. (2024)
Japanese Stable LM
Beta 70B
70
CPT
Llama2 70B
JA,EN
2100 1794 102 stabilityai/japanese-stablelm-base-beta-70b
KARAKURI LM 70B
v0.1
70
CPT
Llama2 70B
JA,EN
2016 1794 18 KARAKURI Inc. (2024)
Llama 2 70B
70
PT
EN
2000 1794 2 Touvron et al. (2023)
Llama 3 70B
70
PT
EN
15000 14250 15 Dubey et al. (2024)
Llama 3 Swallow 70B
70
CPT
Llama3 70B
JA,EN,
Code
15100 14250 115 Fujii et al. (2024)
Swallow 70B
70
CPT
Llama2 70B
JA,EN
2100 1794 102 Fujii et al. (2024)
Qwen2-72B
72
PT
ZH,EN,
Code+27
7000 3500 Yang et al. (2024)
B.2 Evaluation Tasks and Benchmarks